# ]Meta Modern Recommendation System (MRS) RecoWorld: Building Simulated Environments for Agentic Recommender Systems

### v1

@@ 4 Instruction-Following RecSys
Once users are accustomed to giving instructions to LLMs, they tend to do the same with every system they use (Zhou et al., 2023; Zhang et al., 2023; Xu et al., 2025c; Zhuang et al., 2025). For recommender systems, this means developing new agentic capabilities that allow them to respond to user commands and adjust their behavior accordingly. Users might ask to see more or less of the content that interests them, request content from specific sources such as particular friends or topics, or ask questions about the content they are seeing in their screen. This applies to various media forms, ranging from text, photos, videos, multimedia carousels, to products and even social connections.
@@
In the context of recommender systems, the interaction between a user and the system can be modeled as a Markov Decision Process. The task involves a series of states and actions , beginning at an initial state and finally reach the goal state . Here, the environment state represents the user’s mindset, which evolves as the user interacts with recommended items. The state transition is modeled as . It captures how the user’s mindset changes after the recommender system presents a new set of items, and this process can be simulated using a LLM-based user simulator, with parameters . At each step, the agentic recommender system selects an action , yielding a new list of recommendations by considering the user’s explicit or implicit instruction provided by .
@@
The recommender model’s parameters could be related to its perception, memory, planning, tool use modules or the architecture itself in the case of generative recommenders. These parameters are updated iteratively to improve the policy. The ultimate goal is to develop a policy, denoted as , which focuses on identifying the optimal recommendations given the current state , that maximizes the total expected reward over time. A reward function assigns a scalar reward after an action is taken from a given state. We can perform multiple rollouts to estimate cumulative rewards, using on-policy learning methods such as Proximal Policy Optimization (PPO) or off-policy learning techniques such as DPO. The reward function may incorporate metrics such as total session time, number of clicks, self-critique scores, or a combination of factors that measure the system’s ability to follow user instructions and maximize engagement. This approach provides more robust feedback compared to action-level rewards.
@@
Voice-Driven Feedback. There is increasing interest in developing agentic recommender systems. Companies have introduced easy-to-access voice entry points that allow users to adjust the content they see in their feed, friend graph, and across various product surfaces. E.g., a user might say “I want more content about F1,” “Show me more Reels like this,” “Show me what’s new with her,” “Show me content from my high school classmates,” “Show me the latest posts from my family,” “Show me more ‘interesting’ content” and “Show me less political content.” Such user instructions provide valuable feedback signals that agentic recommender systems can learn from.
@@
Instruction-Following (InFoRec) vs. Conversational Recommender System (CRS) Both systems help users discover suitable items through multi-turn interactions. The key difference is that InFoRec respond directly to user commands, while CRS takes a more proactive approach, suggesting items during the conversation. For example, in a CRS chat: User: “Hi, what’s the date today?" CRS: “Today is September 16, 2025. By the way, have you heard of Brad Pitt’s new movie ‘F1’? It’s super popular right now.” (Zhao et al., 2025) Previous research has explored a variety of persuasion strategies, such as authority, emotional appeal, logic, and social proof, to recommend items tailored to users with different personality traits. InFoRec is about following user instructions precisely. It allows a user-facing LLM to manage the conversation, often through voice interactions. Any persuasion strategies are handled by the LLM itself, not by InFoRec. Instead, InFoRec focuses on fine-tuning the list of recommended items by adding, removing, or reordering items among the top results. Think of InFoRec as a tool that is triggered during the conversation: by utilizing the entire conversation and engagement history, InFoRec dynamically adjusts the recommendations shown to the user. This context-aware, instruction-following capability is a crucial step toward developing more agentic recommender systems.
5 Multi-Agent Simulators
@@
Recent advances in multi-agent simulations have substantially enhanced our ability to model the dynamics of individuals making decisions and interacting over time (Park et al., 2023, 2024; Dubois et al., 2024; Gao et al., 2024; Yang et al., 2025; Louie et al., 2025; Qian et al., 2025; Anthis et al., 2025). Particularly, Anthis et al. (2025) highlight the potential of LLM-based social simulations. They identify five key tractable challenges that pave the way for innovations in this space. AgentTorch (Chopra et al., 2024) uses the COVID-19 pandemic to test whether LLM agents can replicate population-level behavioral statistics. It further examines the trade-off between simulation scale and individual agency. OASIS (Yang et al., 2025) simulates social media platforms, e.g., X and Reddit, to study information diffusion, group polarization, and herd behavior. AgentSociety (Piao et al., 2025) provides a testbed for computational social experiments, examining polarization, the spread of inflammatory content, universal basic income, and the impact of shocks such as hurricanes. These simulators vary in scale, from thousands of agents to millions (e.g., OASIS, AgentScope).
@@ 6 Related Work
Evaluation of Recommender Systems. Recommender systems have seen tremendous progress in recent years (Zhai et al., 2024; Wang et al., 2024; Hou et al., 2025; Zhu et al., 2025; eun Yoon et al., 2025; Zhang et al., 2025a; Xia et al., 2025; Deng et al., 2025b; Zhou et al., 2025), with generative recommenders, e.g. Meta’s HSTU and Kuaishou’s OneRec, and emerging agentic recommenders holding promise for delivering highly personalized content. E.g., HSTU achieves up to 65% better ranking metrics than baselines, with a 1.5-trillion-parameter version boosting online A/B metrics by 12.4% (Zhai et al., 2024). Similarly, OneRec replaces the traditional retrieve-then-rank pipeline with an end-to-end generative model and was deployed to millions of users, yielding a +1.6% increase in watch-time. A critical challenge comes with these more powerful models is evaluation. Traditional offline metrics often prove insufficient: promising ideas are dropped due to unconvincing offline results, and improvements that look good offline do not always translate to real user gains. Yet running A/B tests is costly and time-consuming, and it is impractical (and risky) to expose users to every experiment. This has led to growing interest in large-scale simulated environments for recommender evaluation.
@@
Simulated Environments for Training Agentic Systems. Simulated environments have been a driving force in training autonomous agents through RL, ranging from physics sandboxes such as MuJoCo (Zakka et al., 2025) to games like Go and the Arcade Learning Environment (Farebrother and Castro, 2024) to computer-use tasks such as WebArena (Zhou et al., 2024). In the recommender systems domain, one notable effort is KuaiSim, which provides a rich user environment with multi-behavior feedback and cross-session interactions. This enables training RL-based recommenders on sessions and retention scenarios. Another example is AgentRecBench (Shang et al., 2025), which constructs a simulation environment by processing multi-domain datasets (e.g. Yelp, Goodreads, Amazon) and exposing standardized “world” APIs for an agent to query and act. These environments provide a sandbox for agents to plan, explore, and learn complex behaviors without real-world risks, and hold promise to propel the next generation of agentic recommenders.
@@
User Simulation with LLMs. LLM-empowered user simulator has shown promising results in imitating real-world user behaviors across different domains, such as social simulation (Gürcan, 2023), legal judgement (Jiang and Yang, 2025), and recommender systems (Peng et al., 2025). In recommender systems, LLM agents are typically developed to generate simulated user feedback for recommender model evaluation (Yoon et al., 2024) and optimization (Cai et al., 2025). Existing works can be mainly categorized into two groups: 1) Non-conversational user simulator (Zhang et al., 2025b), where the LLM agent mainly give realistic feedback (e.g., like, dislike) on recommendations as exemplified by Agent4Rec (Zhang et al., 2024a) and AgentCF (Zhang et al., 2024b). Typically, non-conversational user simulator cannot proactively give instructions to recommender systems or proactively interact with environments (e.g., other LLM agent). Another line of work lies in 2) conversational user simulator (Guo et al., 2024; Yoon et al., 2024), which endows LLM agents with the flexibility to engage in dialogue with recommender models and other agents (e.g., RecAgent (Wang et al., 2025a)), often incorporating self-reflection mechanisms to model user intent more accurately. To support the development of agentic recommender systems, we borrow the idea from conversational user simulators, which offer a critical advantage in evaluating agentic recommender systems: they can issue context-aware instructions at appropriate moments. For example, indicating a desire to shift away from prior preferences due to fatigue or boredom with previous recommendations.
7 Conclusion
Our RecoWorld offers a novel simulated environment that enables instruction-driven interactions between users and agentic recommender systems. Distinguished from previous works, it features a dual-view architecture, where a simulated user and an agentic recommender engage in multi-turn interactions aimed at maximizing user retention. By modeling realistic user behavior and leveraging multi-turn feedback, RecoWorld provides meaningful reward signals for reinforcement learning agents. It facilitates evaluation using existing datasets and human validation, demonstrating its potential to improve recommendation quality. RecoWorld represents a significant advancement toward more interactive and user-responsive recommender systems in RL settings.

### v2

@@ 4. Instruction-Following RecSys
Once users get used to giving instructions to LLMs, they tend to do the same with every system they use (Zhou et al., 2023; Brooks et al., 2023). For recommender systems, this means developing new agentic capabilities that allow them to respond to user commands and adjust their behavior accordingly. Users might ask to see more or less of the content that interests them, request content from specific sources such as particular friends or topics, or ask questions about the content they are seeing in their screen. This applies to various media forms, ranging from text, photos, videos, multimedia carousels, to products and even social connections.
@@
In the context of recommender systems, the interaction between a user and the system can be modeled as a Markov Decision Process. The task involves a series of interleaved states and actions , beginning at an initial state and finally reach the goal state . Here, the environment state represents the user’s mindset, which evolves as the user interacts with recommended items. The state transition is modeled as . It captures how the user’s mindset changes after the recommender system presents a new set of items, and this process can be simulated using a LLM-based user simulator, with parameters . At each step, the agentic recommender system selects an action , yielding a new list of recommendations by considering the user’s explicit or implicit instruction provided by .
@@
The recommender model’s parameters could be related to its perception, memory, planning, tool use modules or the architecture itself in the case of generative recommenders. These parameters are updated iteratively to improve the policy. The ultimate goal is to develop a policy, denoted as , which focuses on identifying the optimal recommendations given the current state , that maximizes the total expected reward over time. A reward function assigns a scalar reward after an action is taken from a given state. We can perform multiple rollouts to estimate cumulative rewards, using on-policy learning methods such as PPO with asynchronous rewards or off-policy learning such as DPO. The reward function may incorporate metrics such as total session time, number of clicks, self-critique scores, or a combination of factors that measure the system’s ability to follow user instructions and maximize engagement. This approach provides more robust feedback compared to action-level rewards.
@@
Voice-Driven Feedback. There is increasing business interest in developing agentic recommender systems. Companies have introduced easy-to-access voice entry points that allow users to adjust the content they see in their feed, friend graph, and across various product surfaces. E.g., a user might say “I want more content about F1,” “Show me more Reels like this,” “Show me what’s new with her,” “Show me content from my high school classmates,” “Show me the latest posts from my family,” “Show me more ‘interesting’ content” and “Show me less political content.” Such voice-based user instructions provide valuable feedback signals that agentic recommender systems can learn from.
Instruction Alignment. User instructions can span a broad spectrum, so it is necessary to align them accurately with the underlying user intents (Wang et al., 2025b). For example, a user requesting more content on “UFC Fight Night” may have different motivations: they might be looking for a fun local event to enjoy with friends (a short-term, situational interest where other local activities could also appeal), or they could be exploring combat sports, possibly with a long-term interest of training for competitions (in which case Brazilian Jiu-Jitsu would be relevant). Users may provide vague or open-ended feedback, such as “I’m bored,” which poses a greater challenge than responding to specific requests such as “I’d like to see long skirts in light blue for around $200” (Tang et al., 2025b). While lexical or semantic matching can typically align user queries with relevant content, ambiguous instructions require more sophisticated handling. We envision a reasoning-intensive retrieval model powered by LLMs to infer user intents (Xing et al., 2026), refine ambiguous queries, and enhance the system’s ability to understand user instructions.
@@
Instruction-Following (InFoRec) vs. Conversational Recommender System (CRS) The key difference is that InFoRec respond directly to user commands, while CRS takes a more proactive approach, suggesting items during the conversation. For example, in a CRS chat: User: “Hi, what’s the date today?” CRS: “Today is September 16, 2025. By the way, have you heard of Brad Pitt’s new movie ‘F1’? It’s super popular right now” (Wen et al., 2025; Zhao et al., 2025). Previous research has developed various persuasion strategies, such as authority, emotional appeal, social proof, to recommend items tailored to users’ personality traits. In contrast, InFoRec focuses on precisely following user instructions. It refines recommendations by adding, removing, or reordering items among the top results. Think of InFoRec as a tool that is triggered during the conversation: by utilizing the entire conversation and engagement history, InFoRec dynamically adjusts the recommendations shown to the user.
5. Multi-Agent Simulators
@@
Recent advances in multi-agent simulations have substantially enhanced our ability to model the dynamics of individuals making decisions and interacting over time (Park et al., 2023, 2024; Dubois et al., 2024; Gao et al., 2024; Yang et al., 2025; Tang et al., 2025a; Louie et al., 2025; Qian et al., 2025; Anthis et al., 2025). Particularly, Anthis et al. (2025) highlight the potential of LLM-based social simulations. They identify five key tractable challenges that pave the way for innovations in this space. AgentTorch (Chopra et al., 2024) uses the COVID-19 pandemic to test whether LLM agents can replicate population-level behavioral statistics. It further examines the trade-off between simulation scale and individual agency. OASIS (Yang et al., 2025) simulates social media platforms, e.g., X and Reddit, to study information diffusion, group polarization, and herd behavior. AgentSociety (Piao et al., 2025) provides a testbed for computational social experiments, examining polarization, the spread of inflammatory content, universal basic income, and the impact of shocks such as hurricanes. These simulators vary in scale, from thousands of agents to millions (e.g., OASIS, AgentScope).
@@ 6. Related Work
Evaluation of Recommender Systems. Recommender systems have seen tremendous progress in recent years (Wang et al., 2024; Hou et al., 2025; Zhu et al., 2025; Yoon et al., 2025; Xing et al., 2025; Zhang et al., 2025a; Chen et al., 2025c; Deng et al., 2025b; He et al., 2025; Zhou et al., 2026), with generative recommenders, e.g. Meta’s HSTU and Kuaishou’s OneRec, and emerging agentic recommenders holding promise for delivering highly personalized content. For example, HSTU achieves up to 65% better ranking metrics than baselines, with a 1.5-trillion-parameter version boosting online A/B metrics by 12.4% (Zhai et al., 2024). Similarly, OneRec replaces the traditional retrieve-then-rank pipeline with an end-to-end generative model and was deployed to millions of users, yielding a +1.6% increase in watch-time. A critical challenge comes with these more powerful models is evaluation. Traditional offline metrics often prove insufficient: promising ideas are dropped due to unconvincing offline results, and improvements that look good offline do not always translate to real user gains. Yet running A/B tests is costly and time-consuming, and it is impractical (and risky) to expose users to every experiment. This has led to growing interest in large-scale simulated environments for recommender evaluation.
@@
Simulated Environments for Training Agentic Systems. Simulated environments have been pivotal in training autonomous agents through RL, ranging from physics sandboxes such as MuJoCo (Zakka et al., 2025), to games in the Arcade Learning Environment (Farebrother and Castro, 2024), and to computer-use tasks such as WebArena (Zhou et al., 2024). World models, which are learned simulators of environment dynamics, have also played a significant role (Fung et al., 2025). In RecSys, one notable effort is KuaiSim (Zhao et al., 2023), which provides a rich user environment with multi-behavior feedback and cross-session interactions. This enables training RL-based recommenders on sessions and retention scenarios (Zhang et al., 2025b). Another example is AgentRecBench (Shang et al., 2025), which constructs a simulation environment by processing multi-domain datasets (e.g., Yelp, Goodreads, Amazon) and exposing standardized “world” APIs for an agent to query and act. These environments provide a sandbox for agents to plan, explore, and learn complex behaviors.
@@
User Simulation with LLMs. LLM-empowered user simulator has shown promising results in imitating real-world user behaviors across different domains, such as social simulation (Gürcan, 2023), legal judgement (Jiang and Yang, 2025), and recommender systems (Bougie and Watanabe, 2025; Park, 2025; Peng et al., 2025; Chen et al., 2025b; Liu et al., 2026). In RecSys, LLM agents are used to generate simulated user feedback for model evaluation (Yoon et al., 2024; Wei et al., 2025; Chen et al., 2025a; Zhang et al., 2026) and optimization (Cai et al., 2025; Liu et al., 2025). Existing works can be mainly categorized into two groups: 1) Non-conversational user simulators (Zhang et al., 2025c): These LLM agents give realistic feedback (e.g., like, dislike) on recommendations, as exemplified by Agent4Rec (Zhang et al., 2024a) and AgentCF (Zhang et al., 2024b). Typically, non-conversational user simulator cannot proactively give instructions to recommender systems or interact with environments (e.g., other LLM agents). Another line of work lies in 2) Conversational user simulators (Guo et al., 2024; Yoon et al., 2024): These agents engage in dialogue with recommender models and other agents (e.g., RecAgent (Wang et al., 2025a)), often incorporating self-reflection mechanisms to more accurately model user intent. To support the development of agentic RecSys, we borrow the idea from conversational user simulators: they can issue context-aware instructions at appropriate moments. E.g., indicating a desire to shift away from prior preferences due to fatigue or boredom with previous recommendations.
7. Industry Use and Conclusion
Our RecoWorld introduces a novel environment for instruction-driven interactions between users and agentic recommender systems. Internally released in September 2025, this work establishes the foundation for instruction-following recommenders. In this paper, we prioritize presenting a comprehensive framework for building simulated environments. We note that RecoWorld has already received significant interest from leading industrial stakeholders, including Google and Kuaishou RecSys teams, suggesting its potential impact on real-world applications.